# NViST: In the Wild New View Synthesis from a Single Image with Transformers

### v1

@@ Abstract
We propose NViST, a transformer-based model for novel-view synthesis from a single image, trained on a large-scale dataset of in-the-wild images with complex backgrounds. NViST transforms image inputs directly into a radiance field, adopting a scalable transformer-based architecture. In practice, NViST exploits the self-supervised features learnt by a masked autoencoder (MAE), and learns a novel decoder that translates features to 3D tokens via cross-attention and adaptive layer normalization. Our model is efficient at inference since only a single forward-pass is needed to predict a 3D representation, unlike methods that require test-time optimization or sampling multiple times such as 3D-aware diffusion models. We tackle further limitations of current new-view synthesis models. First, unlike most generative models that are trained in a category-specific manner, often on synthetic datasets or on masked inputs, our model is trained on MVImgNet, a large-scale dataset of real-world, casually-captured videos containing hundreds of object categories with diverse backgrounds. Secondly, our model does not require canonicalization of the training data – i.e. aligning all objects with a frontal view – only needing relative pose at training time which removes a substantial barrier to it being used on casually captured datasets. We show results on unseen objects and categories on MVImgNet and even casual phone captures. We conduct qualitative and quantitative evaluations on MVImgNet and ShapeNet to show that our model represents a step forward towards enabling true in-the-wild novel-view synthesis from a single image. The project webpage is https://wbjang.github.io/nvist_webpage.
@@ 1 Introduction
Learning 3D scene representations from RGB images for novel view synthesis or 3D modelling remains a pivotal challenge for the computer vision and graphics communities. Traditional approaches, such as structure from motion (SfM) and multiview stereo pipelines, endeavor to optimize the 3D scene from RGB images directly by leveraging geometric and photometric consistency. The advent of Neural Radiance Fields (NeRF) [43] and its subsequent developments has marked a significant step forward by encoding the entire 3D scene within the weights of a neural network (or feature grid), from RGB images only. Although NeRF requires more than dozens of images for training, efforts have been undertaken to generalize NeRF across multiple scenes by taking a single image as an input [81, 75, 26, 30, 11, 38, 45]. Nonetheless, learning multiple real-world scenes with a NeRF-based model still remains a challenging task as there are scale ambiguities, scene misalignments and diverse backgrounds. The huge success of 2D latent diffusion models [56] has recently sparked interest in 3D diffusion models. One trend is to make diffusion models 3D-aware by fine-tuning a pre-trained 2D diffusion model. However, so far these approaches are trained on masked inputs and do not deal with backgrounds  [41, 42, 40, 52]. Other approaches [23, 9, 69, 13] build the diffusion model on top of volume rendering in 3D space, but these models are computationally expensive and are slow to sample from.
@@
The recent release of large-scale multiview datasets such as Co3D [55] or MVImgNet [82] has meant a huge step forward, which coupled with the availability of robust SfM tools such as COLMAP [62, 61] or ORB-SLAM [47] have enabled the release of their camera parameters. In this paper we exploit a subset of MVImgNet [82] which has one order of magnitude more categories and $\times$2 more objects than Co3D [55].
However, there remain significant challenges to train a new view synthesis model on such large scale datasets due to the huge diversity of objects, scene scales, backgrounds, and scene alignment issues. Motivated by this we propose NViST, a scalable transformer-based architecture trained on a large-scale real-world dataset to enable in-the-wild new view synthesis from a single image. Our contributions can be summed up as follows:
  • Our novel decoder maps MAE features to 3D tokens via cross-attention and adaptive layer normalization conditioned on camera parameters to tackle scale ambiguities.
  • Efficient inference as it only needs one feed-forward pass.
  • NViST only requires relative pose during training and does not require a canonicalized dataset.
  • Our qualitative and quantitative evaluations on test (unseen) scenes and categories show competitive performance on challenging real-world scenes.
  • We demonstrate good performance on a zero-shot new-view synthesis task on out-of-distribution scenes captured on a mobile-phone device.
Figure 2: NViST Architecture. NViST is a feed-forward transformer-based model that takes a single in-the-wild image as input, and renders a novel view. The encoder, a finetuned Masked Autoencoder (MAE), generates self-supervised feature tokens and a class token (cls). Our novel decoder conditioned on normalised focal length and camera distance via adaptive layer normalisation, translates feature tokens to output tokens via cross-attention, and reasons about the occluded region through self-attention. We employ MLPs to reshape the output tokens into a vector-matrix representation that encodes the radiance field and can be queried to obtain the color and sigma of points to perform volume rendering. NViST is trained end-to-end via a combination of losses: photometric $L_{2}$, perceptual $L_{\text{LPIPS}}$, and a distortion-based regulariser $L_{\text{reg}}$. NViST is efficient at inference since it requires only a single forward-pass to predict a 3D from an input image.
@@ 2 Related Work
Transformer and ViT: The Transformer [71], a feed-forward, attention-based architecture, was initially introduced as an alternative to Recurrent Neural Networks. By predicting the next masked tokens with its scalable architecture, it has brought a revolution in the field of natural language processing. Transformer has been used in other domains, including the field of computer vision. Vision Transformer (ViT)  [19] utilizes the patches for attention blocks, and it achieves performance levels comparable to those of Convolutional Neural Networks (CNN)  [24, 65] based architectures in the image recognition task.
Self-supervised learning and MAE: ViT is trained in a supervised way like numerous other vision tasks depending on labels. One of the reasons behind the Transformer’s success story is that it can be trained in a self-supervised way. Masked Autoencoder (MAE) [25] adopts a ViT architecture, and is trained by inpainting the randomly masked patches. MAE undergoes fine-tuning for specific vision tasks, and it demonstrated superior performance to ViT without the need for data augmentation. DINO [6] also adopts the self-supervised learning strategy, and its features are trained on an image level, while MAE is supervised on a patch level.
@@
Neural Implicit Representations: Neural implicit representations aim to learn a 3D representation without direct 3D supervision. It has been explored in previous approaches, including depth prediction [48] and scene representation through ray marching and coordinate-based MLPs [66]. Mildenhall et al. [43] proposed Neural Radiance Fields (NeRF), which integrates coordinate-based MLPs, positional encoding, and volume rendering to perform multi-view stereo of a given scene. Upon optimisation, NeRF can render novel views with impressively high quality. Beyond novel view synthesis, NeRF has found utility in a diverse range of other computer vision tasks such as segmentation [84, 72, 21], surface reconstruction [49, 74, 73, 44], and camera registration [37, 31, 16, 78, 39, 14].
@@
Grid-based representations: One limitation of the original NeRF method is its lengthy training time. As an alternative to using coordinate-based MLPs, grid-based approaches [80, 67, 46, 12, 20, 79] have been proposed to expedite training. In the context of 3D-aware Generative Adversarial Networks (GANs), EG3D [8] introduced the use of triplanes which project 3D features into three different planes. TensoRF [12] compared different tensor decomposition methods and proposed the vector-matrix representation.
@@
Learning multiple scenes based on NeRF: Generalizing NeRFs with multiple scenes has been a challenging problem. Several methods utilise the 2D features from the pre-trained network with the target views [81, 75, 11, 55, 26, 70, 15]. Another approach is to condition the network on latent vectors.  [22, 30, 45, 54, 2, 27]. Adding an adversarial loss to NeRF leads to 3D-aware GAN, allowing consistent rendering from different viewpoints.  [7, 8, 63, 36, 5]. Many of the approaches implicitly assume that the objects are aligned and have the issue of training multiple real scenes.
Diffusion models in 3D: Latent diffusion [56] and the open-source of Stable Diffusion pretrained on the large dataset transform the field of image generation. Directly applying the diffusion models on 3D has issues as there are no large-scale dataset with the 3D ground-truth, and the field of 3D computer vision has evolved into learning 3D only from RGB images. Fine-tuning the pre-trained Stable Diffusion with the large scale synthetic dataset [17] allows the diffusion models to be 3D-aware [41, 52, 40], or the model directly learns to condition on images and camera poses [76]. Single Stage Diffusion NeRF [13] trains the 3D diffusion model and NeRF optimisation together, assuming NeRF intermediate results as the ground-truth. To perform denoising in 3D space, there are approaches [1, 33, 68] that back-project image features to 3D. Instead of denoising in 3D, several methods employ the image-conditioned NeRF to render features in 3D and learn to denoise in 2D spaces  [23, 9, 69]. Diffusion models usually take a long time to sample, and they have the issue of directly modeling the geometry in 3D.
3D Representation with Transformers: Geometry-free methods employing Transformer architectures have been explored as seen in  [35, 57, 58, 59]. VisionNeRF [38] utilizes the ViT encoder to extract features, which are further mapped into multi-level feature maps via CNN decoder. VQ3D [60] extends the vector quantization models for 3D, and successfully trains ImageNet dataset with the pseudo disparity and adversarial losses. MCC [77] takes a single RGB-D as an input, and predicts the color, occupancy. GINA-3D [64] leverages cross-attention between 2D patches and triplanes, and and focuses more on quantization and category-specific generation. Shap-E [32] is based on diffusion models, and it generates 3D implicit functions from point clouds and images. The concurrent work LRM [28] extracts image features through DINO [6], and refines learnable positional embeddings via attention mechanisms. However, LRM trains the dataset without backgrounds, deploys a larger model (more than $500$M parameters) and requires much more computing ($128$ A100 GPUs), while NViST learns scenes with backgrounds, and is trained with a single A100 GPU.
@@ 3 Methodology
Figure 3: Encoder Output Visualisation: (Top) Input images. (Middle) Features from a fine-tuned MAE, which serves as initialisation to our encoder. (Bottom) Feature tokens after end-to-end training, visualized after dimensionality reduction with PCA to $3$ dimensions. Features appear smoother and more segment-focused, supporting the fact that updating encoder weights significantly improves the performance of the model, and is further evidenced by the quantitative analysis in Table 1
NViST is a feed-forward conditional encoder-decoder model built upon transformers to predict a radiance field from a single image to enable synthesis from novel viewpoints. As Figure 2 shows, our architecture is structured into three key components: two transformer-based modules (encoder and decoder) and a NeRF renderer, which is composed of a shallow multi-layer perceptron (MLP) and a differentiable volume rendering module. NViST has approximately $216$M parameters: $85$M for the encoder, $131$M for the decoder, and $7$K for the renderer.
@@ 3.1 Encoder
Input images are first split into a set of fixed size non-overlapping patches before being fed to the encoder, which predicts feature tokens using a ViT-B transformer architecture. We formulate the encoder $\mathcal{E}$ as $F,C=\mathcal{E}(I)$ where $I$ are input images and $F$ and $C$ are the feature and class tokens respectively. In practice, we use the encoder of a pre-trained MAE [25], a self-supervised vision learner trained to predict masked image patches, which we further finetune on the training dataset to adapt to the different image resolution from ImageNet, on which MAE [25] is trained. As illustrated in Figure  3, we find that the self-supervised features learnt by the MAE [25] encapsulate a powerful intermediate representation of the scene’s geometric and appearance properties that can be leveraged by the decoder to predict the radiance field. Note that the weights of the encoder are continuously updated during end-to-end training, since we found that this enhances the encoder’s ability to generate smoother and more segment-focused features, as in Figure  3. This leads to the better performance of the model as shown in Table 1.
@@ 3.2 Decoder
The decoder $\mathcal{D}$ has output tokens $O$ as learnable parameters. For simplicity, from here on we will use the term output tokens $O$ to refer to the concatenation of output tokens and class token $C$, to which we further add positional embeddings similarly to [71].
@@ 3.2.1 Relative Camera Pose
A key strength of our approach is that it does not require all objects/scenes in the dataset to be brought into alignment by expressing their pose with respect to a canonical reference frame. Instead, we assume the rotation of the camera associated with the input image $I_{i}$ to be the identity and we express the rotation of any other image $I_{j}$ as a relative rotation $R_{i\rightarrow j}$. We assume that the camera location of the input image is $T_{i\rightarrow i}=(0,0,z)$, where $z$ is the normalized distance of the input camera from the origin of the world coordinate frame, located at the centroid of the 3D bounding box containing the point cloud, and $T_{i\rightarrow j}={R_{i}}^{T}(T_{j}-T_{i})$.
@@ 3.2.2 Conditioning on Camera Parameters
There are several ways to apply conditioning such as concatenating camera parameters to output tokens $O$ or feature tokens $F$. However, in NViST we apply adaptive layer normalization, as the camera parameters influence the overall scale of the scene. Conditioning camera parameters improves the model performance as in Table 1.
@@ 3.2.3 Attention Blocks
Finally output tokens are updated via residual connection $O\leftarrow O+\gamma\times\text{Attn}\cdot F_{n}$, where $\gamma$ is the gate scale also regressed by $\mathcal{A}$. Note that self-attention is obtained in an equivalent way, just between output tokens.
@@ 3.3 Rendering
While the value of density $\mathbf{\sigma}$ is obtained by applying ReLU activation directly to the feature value $\mathcal{T}_{\mathbf{x}}$ at point $\mathbf{x}$, the colour $\mathbf{c}$ is predicted with a shallow MLP, conditioned on the viewing direction. The VM representation outperforms the method of aggregating features by projecting them onto three 2D planes and then concatenating them as in Table 1.
@@ 3.4 Training Losses
We employ a combination of losses to train our architecture in an end-to-end manner including the L2 photometric rendering loss, a Learned Perceptual Image Patch Similarity (LPIPS) loss [83], and the distortion-based regulariser proposed by  [3].
Given ground-truth pixel colors $\bf{v}$, estimates $\bf{\hat{v}}$ and accumulated transmittance values $\bf{w}$ along the points $\bf{x}$ on the ray, our overall loss function is defined as
@@
For MVImgNet we set $\lambda=0.1$ and $\beta=0.001$. We found that LPIPS to be extremely effective for real world datasets as in Table 1.

### v2

@@ Abstract
We propose NViST, a transformer-based model for efficient and generalizable novel-view synthesis from a single image for real-world scenes. In contrast to many methods that are trained on synthetic data, object-centred scenarios, or in a category-specific manner, NViST is trained on MVImgNet, a large-scale dataset of casually-captured real-world videos of hundreds of object categories with diverse backgrounds. NViST transforms image inputs directly into a radiance field, conditioned on camera parameters via adaptive layer normalisation. In practice, NViST exploits fine-tuned masked autoencoder (MAE) features and translates them to 3D output tokens via cross-attention, while addressing occlusions with self-attention. To move away from object-centred datasets and enable full scene synthesis, NViST adopts a 6-DOF camera pose model and only requires relative pose, dropping the need for canonicalization of the training data, which removes a substantial barrier to it being used on casually captured datasets. We show results on unseen objects and categories from MVImgNet and even generalization to casual phone captures. We conduct qualitative and quantitative evaluations on MVImgNet and ShapeNet to show that our model represents a step forward towards enabling true in-the-wild generalizable novel-view synthesis from a single image. Project webpage: https://wbjang.github.io/nvist_webpage.
@@ 1 Introduction
Learning 3D scene representations from RGB images for novel view synthesis or 3D modelling remains a pivotal challenge for the computer vision and graphics communities. Traditional approaches, such as structure from motion (SfM) and multiview stereo pipelines, endeavor to optimize the 3D scene from RGB images directly by leveraging geometric and photometric consistency. The advent of Neural Radiance Fields (NeRF) [52] and its subsequent developments has marked a significant step forward by encoding the entire 3D scene within the weights of a neural network (or feature grid), from RGB images only. Although NeRF requires more than dozens of images for training, efforts have been undertaken to generalize NeRF across multiple scenes by taking a single image as an input [99, 91, 31, 36, 15, 46, 54, 63]. Nonetheless, generalizing NeRF-based models to multiple real-world scenes remains a challenge due to scale ambiguities, scene misalignments and diverse backgrounds. The huge success of 2D latent diffusion models [66] has sparked interest in 3D diffusion models. One trend is to make diffusion models 3D-aware by fine-tuning a pre-trained 2D diffusion model. However, so far these approaches are trained on centered objects, masked inputs, do not deal with backgrounds, and assume a simplified camera model (3-DOF) [49, 48, 61, 76]. Other approaches [28, 13, 84, 17] build the diffusion model on top of volume rendering in 3D space, but they are computationally expensive and slow to sample from.
@@
The recent release of real-world large-scale multiview datasets such as Co3D [65] or MVImgNet [100], coupled with the availability of robust SfM tools such as COLMAP [72, 71] or ORB-SLAM [56] to enable camera pose estimation, has opened the door to large-scale training for new view synthesis. However, significant challenges remain to train a scene-level new view synthesis model on such real-world, large scale datasets due to the huge diversity of objects, categories, scene scales, backgrounds, and scene alignment issues. Motivated by this we propose NViST, a transformer-based architecture trained on a large-scale dataset to enable in-the-wild new view synthesis (NVS) from a single image. We exploit a subset of MVImgNet [100] which has one order of magnitude more categories and $\times$2 more objects than Co3D [65]. Our contributions can be summed up as follows:
  • NViST can model general real-world scenes, including backgrounds, only requiring relative pose during training.
  • Our novel decoder maps MAE features to 3D output tokens via cross-attention, conditions on camera parameters via adaptive layer normalisation and addresses occlusions with self-attention.
  • Our qualitative and quantitative evaluations on MVImgNet test sequences show good performance on challenging real-world scenes.
  • We demonstrate good generalization results on a zero-shot new-view synthesis task, on phone-captured scenes.
Figure 2: Architecture. NViST is a feed-forward transformer-based model that takes a single in-the-wild image as input, and renders a novel view. The encoder, a finetuned Masked Autoencoder (MAE), generates feature tokens, which are translated to output tokens via cross-attention by our novel decoder, conditioned on normalised focal length and camera distance via adaptive layer normalisation. Self-attention blocks allow reasoning about occlusions. Output tokens are reshaped into a vector-matrix representation that is used for volume rendering. NViST is trained end-to-end via a balance of losses: photometric $L_{2}$, perceptual $L_{\text{LPIPS}}$, and a distortion-based regulariser $L_{\text{reg}}$.
@@ 2 Related Work
Transformer, ViT and MAE: The transformer [86], a feed-forward, scalable, attention-based architecture, has brought a revolution to the field of natural language understanding. Inspired by it, the vision transformer (ViT)  [24] uses image patches as tokens to achieve performance levels comparable to those of CNNs [29, 78] in many computer vision tasks. While the ViT is trained in a supervised way, masked autoencoders (MAE) [30] can be trained in a self-supervised way by randomly masking and in-painting patches, and be further fine-tuned on specific tasks.
@@
Neural Implicit Representations: Neural implicit representations aim to learn a 3D representation without direct 3D supervision using neural networks. They have been employed for various tasks, including depth prediction [57] and scene representation through ray marching and coordinate-based MLPs [79]. Mildenhall et al. [52] proposed Neural Radiance Fields (NeRF), which integrates coordinate-based MLPs, positional encoding, and volume rendering to encode a scene in the weights of neural network. Upon optimisation, novel views can be rendered with impressively high quality. Beyond novel view synthesis, NeRF has found utility in a diverse range of computer vision tasks such as segmentation [102, 87, 26, 10], surface reconstruction [58, 90, 89, 53, 7], and camera registration [45, 37, 20, 96, 47, 18, 92, 50].
@@
Grid-based representations: A limitation of the original NeRF method is its lengthy training time. As an alternative to coordinate-based MLPs, grid-based approaches [98, 81, 55, 16, 25, 97] have been proposed to expedite training. TensoRF [16] proposed the vector-matrix (VM) representation as an efficient and compact way to represent the 3D grid. In the context of 3D-aware Generative Adversarial Networks (GANs), EG3D [12] introduced triplanes by projecting 3D features into three different planes. Several approaches use the triplane representation to learn to generate 3d implicit representations from ImageNet [80, 70, 74, 64].
@@
Learning multiple scenes using NeRF: Generalising NeRFs to multiple scenes remains a challenging problem. Several methods associate 2D features with the target views [99, 91, 15, 65, 31, 85, 19, 35], while others condition the network on latent vectors with a shared MLP across the dataset [27, 36, 54, 63, 3, 32]. Adding an adversarial loss to NeRF leads to 3D-aware GANs, which allow consistent rendering from different viewpoints [11, 12, 73, 44, 8]. Most approaches that use a single input image require aligned datasets like ShapeNet [14, 79] or FFHQ [40], and fail to deal with scale ambiguities or diverse backgrounds.
Diffusion for NVS: Latent diffusion [66] and its open-source release Stable Diffusion have transformed the field of image generation. However, applying diffusion models to learn 3D implicit representations is not straightforward as there is no access to 3D ground-truth. 3DiM [94] proposes a pose-conditioned image-to-image approach. Dreamfusion [60] introduced score distillation sampling (SDS) to train NeRF with a 2D diffusion model. SDS has been used in many follow-up works [51, 77, 83, 88, 93]. Fine-tuning Stable Diffusion on a large-scale synthetic dataset [21] allows diffusion models to be 3D-aware such as Zero-1-to-3 and its follow-ups [49, 61, 48, 76]. Other approaches denoise directly in 3D and supervise the model in 2D space after rendering,  [39, 82, 2, 28, 13, 84, 1], use the 2D diffusion model as a prior [22], or optimise NeRF jointly and regard it as the ground truth [17]. Similar to latent diffusion [41, 74, 5] adopt 2-stage training.
3D Representation with Transformers: Geometry-free methods employing transformer architectures have been explored as seen in  [43, 67, 68, 69]. Others build NeRF representations using transformers [46, 70, 75, 95, 38]. The concurrent work LRM [33] extracts image features through DINO [9], and refines learnable positional embeddings via attention mechanisms. Unlike  NViST, LRM focuses on object-centric scenes without background and employs triplanes instead of vector-matrix representation.
@@ 3 Methodology
NViST is a feed-forward conditional encoder-decoder model built upon transformers to predict a radiance field from a single image to enable synthesis from novel viewpoints. As Figure 2 shows, our architecture is structured into three key components: two transformer-based modules (encoder and decoder) and a NeRF renderer, which is composed of a shallow multi-layer perceptron (MLP) and a differentiable volume rendering module.
@@ 3.1 Encoder
Input images are first split into a set of fixed size non-overlapping patches before being fed to the encoder, which predicts feature tokens using a ViT-B transformer architecture. We formulate the encoder $\mathcal{E}$ as $F,C=\mathcal{E}(I)$ where $I$ are input images and $F$ and $C$ are the feature and class tokens respectively. In practice, we use the encoder of a pre-trained MAE [30], a self-supervised vision learner trained to predict masked image patches, which we further finetune on the training dataset to adapt to the different image resolution from ImageNet, on which MAE [30] is trained. As illustrated in Figure  3, we find that the self-supervised features learnt by the MAE [30] encapsulate a powerful intermediate representation of the scene’s geometric and appearance properties that can be leveraged by the decoder to predict the radiance field. Note that the weights of the encoder are continuously updated during end-to-end training, since we found that this enhances the encoder’s ability to generate smoother and more segment-focused features, as shown in Figure  3, and better performance as shown in Table 2.
Figure 3: Encoder Output Visualisation: (Top) Input images. (Middle) Features from a fine-tuned MAE, which serve as initialisation to our encoder. (Bottom) Features after end-to-end training. Features shown after reducing to $3$ dimensions with PCA. Optimised features appear smoother and more segment-focused, supporting the fact that updating encoder weights significantly improves the performance (see also ablation in Table 2).
@@ 3.2 Decoder
The decoder $\mathcal{D}$ has output tokens $O$, which are learnable parameters, initialised following a random normal distribution. For simplicity, from here on we will use the term output tokens $O$ to refer to the concatenation of output tokens and class token $C$, to which we further add positional embeddings [86].
@@ 3.2.1 Relative Camera Pose
A key strength of our approach is that it does not require all objects/scenes in the dataset to be brought into alignment by expressing their pose with respect to a canonical reference frame. Instead, we assume the rotation of the camera associated with the input image $I_{i}$ to be the identity and we express the rotation of any other image $I_{j}$ as a relative rotation $R_{i\rightarrow j}$. We assume that the camera location of the input image is $T_{i\rightarrow i}=(0,0,z)$, where $z$ is the normalized distance of the input camera from the origin of the world coordinate frame, located at the centroid of the 3D bounding box containing the point cloud, and $T_{i\rightarrow j}={R_{i}}^{T}(T_{j}-T_{i})$. Zero-1-to-3 [49] also adopts a relative-pose based approach, however it assumes objects are located in the centre and uses a 3-DoF camera pose model (radius, elevation and azimuth), while NViST uses a full 6-DoF model for the camera pose.
@@ 3.2.2 Conditioning on Camera Parameters
There are several ways to apply conditioning such as concatenating camera parameters to output tokens $O$ or feature tokens $F$. However, in NViST we apply adaptive layer normalization, as the camera parameters influence the overall scale of the scene. Conditioning camera parameters improves the model performance as seen in Table 2.
@@
An alternative to adaptive layer normalisation would be to concatenate $M$ to output tokens $O$, but as seen in our ablation (Table 2) this strategy does not lead to better results.
@@ 3.2.3 Attention Blocks
Finally, output tokens are updated via residual connection $O\leftarrow O+\gamma\times\text{Attn}\cdot F_{n}$, where $\gamma$ is the gate scale also regressed by $\mathcal{A}$. Note that self-attention is obtained in an equivalent way, just between output tokens.
@@ 3.3 Rendering
While the value of density $\mathbf{\sigma}$ is obtained by applying ReLU activation directly to the feature value $\mathcal{T}_{\mathbf{x}}$ at point $\mathbf{x}$, the colour $\mathbf{c}$ is predicted with a shallow MLP, conditioned on the viewing direction. The VM representation outperforms using a triplane as shown in our ablation study (Table 2).
@@ 3.4 Training Losses
We employ a combination of losses to train our architecture in an end-to-end manner including the L2 photometric rendering loss, a Learned Perceptual Image Patch Similarity (LPIPS) loss [101], and the distortion-based regulariser proposed by  [4]. Given ground-truth pixel colors $\bf{v}$, estimates $\bf{\hat{v}}$ and accumulated transmittance values $\bf{w}$ along the points $\bf{x}$ on the ray, our overall loss function is defined as
@@
For MVImgNet $\lambda=0.1$ and $\beta=0.01$. We found LPIPS to be extremely effective on real world datasets (see Table 2).